OpenStudy (anonymous):

Solve log4 x – 1 = 2

OpenStudy (anonymous):

change it to exponential 4^2=x-1 ,and then solve

OpenStudy (anonymous):

this all depends how this is written \[\log_{4} (x-1)=2?\]

OpenStudy (anonymous):

\[\log_{4}(x)-1=2? \]